### v1

# Learning Survival Models with Right-Censored Reporting Delays

Abstract
Survival analysis is a statistical technique used to estimate the time until an event occurs. Although it is applied across a wide range of fields, adjusting for reporting delays under practical constraints remains a significant challenge in the insurance industry. Such delays render event occurrences unobservable when their reports are subject to right censoring. This issue becomes particularly critical when estimating hazard rates for newly enrolled cohorts with limited follow-up due to administrative censoring. Our study addresses this challenge by jointly modeling the parametric hazard functions of event occurrences and report timings. The joint probability distribution is marginalized over the latent event occurrence status. We construct an estimator for the proposed survival model and establish its asymptotic consistency. Furthermore, we develop an expectation-maximization algorithm to compute its estimates. Using these findings, we propose a two-stage estimation procedure based on a parametric proportional hazards model to evaluate observations subject to administrative censoring. Experimental results demonstrate that our method effectively improves the timeliness of risk evaluation for newly enrolled cohorts.
Figure 1: Overview of problem setting. There are two processes: one leading up to an accident and another from the accident to its corresponding report. Without a report, the accident information enclosed in parentheses is not observed. The main objective of this study is to estimate the accident hazard function for the timely risk evaluation of a newly enrolled cohort with homogeneous risk. Given the limited observation window for such cohorts, the parameter estimator in the survival model is required to achieve asymptotic consistency even under administrative censoring.
1 Introduction
Survival analysis (Kalbfleisch and Prentice 2002) is a statistical method for estimating the time until the occurrence of an event. Its applications span a wide range of fields, including public health, biomedicine, reliability engineering, finance, and marketing. Within this framework, many approaches have been developed to flexibly capture various types of observations (Wang, Li, and Reddy 2019). The Cox proportional hazards model (Cox 1972) is widely used to represent time dependence and covariate effects, assuming a constant hazard ratio over time. Its hazard function is written as the product of a non-parametric baseline hazard and a covariate-dependent risk component. The regression coefficients of the covariates are estimated via partial likelihood (Cox 1975), without specifying the baseline hazard. After the coefficients are obtained, the baseline hazard is estimated (Breslow 1972). In subsequent works, the baseline hazard function has been parametrically modeled using piecewise constant functions (Holford 1980) or B-splines (Royston and Parmar 2002). This parameterization enables simultaneous full likelihood-based estimation of both time dependence and covariate effects.
In practical applications across certain industries, events are not reported immediately but with a delay, typically due to operational or infrastructural constraints. Applying survival analysis techniques in such settings presents a significant challenge, as event occurrences remain unobserved until they are reported after a delay. Ignoring these delays can lead to an underestimation of event frequencies. The epidemiological literature highlights the importance of adjusting for reporting delays. In typical clinical trials, there is a time lag between the occurrence of an event and the ascertainment of vital status. To account for this delay, alternative estimators to the naive Kaplan-Meier have been proposed in previous studies (Hu and Tsiatis 1996; der Laan and Hubbard 1998). In infectious disease research, it is often difficult to clearly define the population under observation, leading to the truncation of individuals who remain asymptomatic or go unreported. Existing research (Lagakos, Barraj, and Gruttola 1988; Kalbfleisch and Lawless 1991; Pagano et al. 1994) has estimated delay distributions using truncated data.
One of the key applications of modeling reporting delays lies in insurance risk management, which is essential for maintaining the long-term stability of insurance companies and ensuring fairness among stakeholders. As outlined in section 2, a variety of approaches have been proposed for reserving incurred but not reported (IBNR) claims. Since reserving typically focuses on portfolio-level estimates, these approaches do not directly support risk evaluation at the individual policy level. Additionally, unlike truncated data, insurance policies are typically subject to right censoring. Incorporating censored individuals who are not entirely absent but only partially observed has not yet been well established. As a result, the practical analysis of individual risk profiles often requires waiting until there is sufficient confidence that most accidents have been reported. This waiting period does not pose a significant problem when underwriting policies within homogeneous and historically well-understood risk segments. However, untimely risk evaluation for new segments with high uncertainty may critically affect the financial position of insurance companies. This also implies that timely decision-making such as repricing or product discontinuation can help broaden the range of acceptable risk profiles. Consequently, individual risk evaluation that accounts for right-censored reporting delays is highly appealing in efforts to expand the reach of insurance value.
In this study, we explore a survival analysis framework that accounts for right-censored reporting delays. Figure 1 provides an overview of our problem setting. To capture the structure induced by reporting delays, we model unreported event occurrences as latent variables and construct the joint probability distribution marginalized over them. The corresponding estimator based on this marginal distribution is shown to be consistent, and its estimates can be computed using the expectation-maximization algorithm. However, when evaluating the risk of a newly enrolled cohort in a timely manner, the observation period is often limited due to administrative censoring. This limitation makes it difficult to ensure the consistency of the parameter estimator. To address this issue, we propose a two-stage estimation procedure based on a parametric proportional hazards model, illustrated in fig. 2. The proposed procedure yields a consistent estimator in the target domain by transferring parameters from the source domain without administrative censoring. Experimental results demonstrate that our approach enhances the timeliness of insurance risk evaluation.
Organization. The remainder of this paper is organized as follows: Section 2 reviews existing research on IBNR claim estimation. Section 3 covers preliminaries necessary for introducing our approach. Section 4 defines the problem setting. Section 5 describes our methodology. Section 6 presents the experimental results. Section 7 concludes this study. Appendix A provides a summary of the notations introduced in the main text. Appendix B includes supplementary material on the preliminaries. Appendix C provides the proofs of the theoretical results.
Figure 2: Overview of the two-stage estimation procedure. The target domain, representing a newly enrolled cohort, is subject to administrative censoring. Due to identifiability limitations inherent in survival models, the covariate effect in the target domain cannot be estimated without supplementary information. To address this challenge, we consider a setting in which supplementary data from a source domain unaffected by administrative censoring are available.
2 Related Work
Reserving for IBNR claims is a fundamental topic in actuarial science (Skurnick 1973; Taylor 2000). The chain ladder method (Mack 1993, 1999) and the Bornhuetter-Ferguson method (Bornhuetter and Ferguson 1972; Mack 2008) are widely used in practice. These methods estimate reporting delay structures using two-dimensional tables that aggregate historical claim amounts by accident and reporting dates. In recent decades, a range of more flexible models has been developed to address the limitations of these traditional techniques, particularly by capturing individual-level claim arrival processes (Antonio and Plat 2014) and incorporating environmental variation (Badescu, Lin, and Tang 2016; Badescu et al. 2019). Verbelen et al. (2022) integrated the processes of event occurrences and reporting delays using latent variables at the aggregate level.
Existing approaches have been developed to improve the accuracy of reserving at the portfolio level. In contrast, we focus on timely risk evaluation at the individual policy level, with applications to pricing and underwriting. Note that the joint parameter estimation used in the first stage of the two-stage procedure can also be applied to reserving tasks.
3 Preliminaries
Throughout this paper, we denote random variables by uppercase letters and their realizations by lowercase letters, omitting indices and accents such as in $\bar{Y}$ and $y_{i}$. We assume that all probability density functions are positive, bounded, and continuous. Furthermore, the parameter spaces associated with these functions are assumed to be compact and to contain the true parameter values. A prime attached to a parameter signifies that it belongs to the same parameter space, as in $\bm{\theta}$ and $\bm{\theta}^{\prime}$. A semicolon is used in the notation to explicitly specify parameters, as in $f(y;\bm{\theta})$. A vertical bar indicates conditioning on observed covariates, as in $f(y\mid\bm{x})$. Let $\mathbb{I}$ denote the indicator function.
Survival analysis models the time until an event occurs as a non-negative random variable $T$. The hazard function of $T$ at time $t\in[0,\infty)$ is defined as $h(t)\equiv f(t)S(t)^{-1}$, where $f(t)$ denotes the probability density function of $T$, and the corresponding survival function is given by $S(t)\equiv 1-\int_{0}^{t}f(s)ds$. The censoring time is modeled as a non-negative random variable $C$, assumed to be independent of $T$. Define the observation time as $Y\equiv\min\{T,C\}$ and the event indicator as $V\equiv\mathbb{I}(T\leq C)$. For $y\in[0,\infty)$ and $v\in\{0,1\}$, the joint distribution of $(Y,V)$ is
$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1),\end{cases}$ (1)
where $f_{c}$ and $S_{c}$ denote the probability density function and the survival probability of $C$, respectively. Administrative censoring refers to situations where observations are censored at a predetermined time point $\tau\in[0,\infty)$. Let $\bar{Y}\equiv\min\{Y,\tau\}$ and $\bar{V}\equiv\mathbb{I}(T\leq C,T\leq\tau)$. For $y\in[0,\tau]$ and $v\in\{0,1\}$, the joint distribution of $(\bar{Y},\bar{V})$ is
$\bar{g}(y,v)\equiv\begin{cases}S(\tau)S_{c}(\tau)&(v=0,y=\tau)\\ S(y)f_{c}(y)&(v=0,y\neq\tau)\\ f(y)S_{c}(y)&(v=1).\end{cases}$ (2)
The hazard functions of $T$ and $C$ are learned from independent observations $\mathcal{D}\equiv(\bm{x}_{i},y_{i},v_{i})_{i=1}^{n}$. Each observation is characterized by the covariates $\bm{x}_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$, the time $y_{i}$ until the event or censoring, and the event indicator $v_{i}\in\{0,1\}$, where $y_{i}\in[0,\infty)$ in the absence of administrative censoring, $y_{i}\in[0,\tau]$ otherwise, and the event indicator takes $1$ when the event occurs before censoring. For simplicity, the realizations of $(Y,V)$ and $(\bar{Y},\bar{V})$ are not distinguished, and are collectively denoted by $(y_{i},v_{i})$.
The hazard functions of $T$ and $C$ are parameterized by unknown real-valued vectors $\bm{\theta}$ and $\bm{\theta}_{c}$, respectively. A simple example of a parametric hazard function is a constant hazard rate defined by a single parameter $\lambda\in(0,\infty)$. The log likelihood is given by $\mathcal{L}(\bm{\theta})+\mathcal{L}_{c}(\bm{\theta}_{c})$, where the two components are defined as
$\mathcal{L}(\bm{\theta})\equiv\log\prod_{i=1}^{n}f(y_{i}\mid\bm{x}_{i})^{v_{i}}S(y_{i}\mid\bm{x}_{i})^{1-v_{i}},$ (3)
$\mathcal{L}_{c}(\bm{\theta}_{c})\equiv\log\prod_{i=1}^{n}S_{c}(y_{i}\mid\bm{x}_{i})^{v_{i}}f_{c}(y_{i}\mid\bm{x}_{i})^{1-v_{i}}.$ (4)
Based on the discrepancy between eq. 1 and eq. 2, we derive the corresponding likelihood function under administrative censoring. As an alternative to $\mathcal{L}_{c}$, we define $\mathcal{L}_{c}^{\dagger}$ as
$\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})\equiv\log\prod_{i=1}^{n}S_{c}(y_{i}\mid\bm{x}_{i})^{v_{i}^{\dagger}}f_{c}(y_{i}\mid\bm{x}_{i})^{1-v_{i}^{\dagger}},$ (5)
where $v_{i}^{\dagger}\equiv\mathbb{I}(y_{i}\geq\tau(1-v_{i}))$. Then, the log likelihood is given by $\mathcal{L}(\bm{\theta})+\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})$. The estimator of $\bm{\theta}$ is obtained by maximizing $\mathcal{L}(\bm{\theta})$, regardless of whether administrative censoring is present. A probability density function is said to be identifiable if the same distribution corresponds to only one parameter value. Provided that this condition holds for $f$, the consistency of the estimator is established in section B.1 and section B.2. This property is useful when the primary focus is on the hazard function of $T$, and that of $C$ is regarded as a nuisance.
Figure 3: Notations for observations. The figure illustrates three possible patterns of accident and report status under a common right censoring time. In the top two patterns where $v_{i}=0$, the accident status $(z_{i},w_{i})$ remains unobserved.
4 Problem Setting
In this study, we consider an unconventional setting in which two events occur sequentially over time, and jointly model both the time to the earlier event and the time from it to the later one. The earlier event corresponds to an accident, while the later event represents its report. We attach subscripts $1$ and $2$ to the symbols $(T,h,f,S,\bm{\theta})$ to refer to the earlier and later events, respectively. Assume that $T_{1},T_{2}$, and $C$ are mutually independent. Unless otherwise stated, $\bm{\theta}_{1},\bm{\theta}_{2}$, and $\bm{\theta}_{c}$ are assumed to be unknown. Let $Z\equiv\min\{T_{1},C\},W\equiv\mathbb{I}(T_{1}\leq C),\bar{Z}\equiv\min\{Z,\tau\}$, and $\bar{W}\equiv\mathbb{I}(T_{1}\leq C,T_{1}\leq\tau)$. Define $Y,V,\bar{Y}$, and $\bar{V}$ by substituting $T=T_{1}+T_{2}$ into their respective expressions. Each observation is characterized by the covariates $\bm{x}_{i}\in\mathcal{X}$, the time $z_{i}\in[0,y_{i}]$ until the earlier event or censoring, the event indicator $w_{i}\in\{0,1\}$ for the earlier event, the time $y_{i}$ until the later event or censoring, and the event indicator $v_{i}$ for the later event. For simplicity, the same notations are used to represent the realizations of both $(Z,Y,W,V)$ and $(\bar{Z},\bar{Y},\bar{W},\bar{V})$.
A similar setting involving two successive events under right censoring is also discussed by Visser (1996). We additionally consider the setting in which the earlier event is unobserved until the later event occurs. The triplet $(\bm{x}_{i},y_{i},v_{i})$ is observed for each $1\leq i\leq n$, while the pair $(z_{i},w_{i})$ is observed only if $v_{i}=1$. The probability of observing $(w,v)=(0,1)$ is zero by the definitions of $(W,V)$ and $(\bar{W},\bar{V})$. Figure 3 illustrates three possible patterns that arise under this extended setting. For simplicity, we set $v_{i}=0$ for $1\leq i\leq m$ and $v_{i}=1$ for $m+1\leq i\leq n$. We also assume that $\sum_{i=m+1}^{n}z_{i}>0$.
The main objective is to evaluate the risk associated with a newly enrolled cohort that exhibits homogeneous risk characteristics. When this evaluation is conducted in a timely manner, each individual’s follow-up duration from enrollment to the evaluation date is bounded above by a fixed time limit $\tau$, which is the administrative censoring. Under the absence of administrative censoring, the consistency of the parameter estimator can be shown under a moderate assumption, as described in theorem 5.3. However, in the presence of administrative censoring, proving consistency requires too restrictive assumptions on the survival models in our problem setting. To address this limitation, we use additional observations from a source domain that is not subject to administrative censoring. Additionally, we assume a transfer condition stating that the time dependence of accident occurrences and the delay distribution in the source domain are the same as those in the target domain for the newly enrolled cohort. Since insurance companies typically maintain long-term data on diverse policyholders, it is reasonable to assume the existence of comparable cohorts unaffected by administrative censoring.
In these settings, we propose a two-stage estimation procedure. In the second stage, applied to the target domain, we construct a consistent estimator under administrative censoring by leveraging information from the source domain and the transfer condition. The details of the two-stage estimation procedure are provided in the subsequent section.
5 Approach
In section 5.1, we develop parameter estimators that account for right-censored reporting delays under both the presence and absence of administrative censoring. In section 5.2, we propose an approach for jointly estimating the parameters of the earlier and later hazard functions under the absence of administrative censoring. These methods constitute the two-stage estimation procedure presented in section 5.3.
5.1 Estimator Formulation
We first construct the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ in the absence of administrative censoring. The corresponding probability model can be derived through a straightforward calculation.
Lemma 5.1.
For $0\leq y$ and $v=0$, the joint distribution of $(Y,V)$ is
$g_{v=0}(y)\equiv S_{\circ}(y)f_{c}(y),$ (6)
where $S_{\circ}(y)\equiv S_{1}(y)+\int_{0}^{y}f_{1}(t)S_{2}(y-t)dt$. For $0\leq z\leq y$ and $v=1$, the joint distribution of $(Z,Y,V)$ is
$g_{v=1}(z,y)\equiv f_{\circ}(z,y)S_{c}(y),$ (7)
where $f_{\circ}(z,y)\equiv f_{1}(z)f_{2}(y-z)$.
Applying lemma 5.1, the log likelihood can be decomposed into $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}_{c}(\bm{\theta}_{c})$, where $\mathcal{E}$ is defined by
$\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\log\prod_{i=1}^{n}S_{\circ}(y_{i}\mid\bm{x}_{i})^{1-v_{i}}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})^{v_{i}}.$ (8)
To estimate $(\bm{\theta}_{1},\bm{\theta}_{2})$, we disregard the second term of the log likelihood and maximize the first. Although the structure of the joint distribution resembles that of eq. 1, it involves a newly introduced pair of functions, $(S_{\circ},f_{\circ})$. In this case, the identifiability of $(g_{v=0},g_{v=1})$ with respect to $(\bm{\theta}_{1},\bm{\theta}_{2})$ is not straightforward, since the realization of $(Z,W)$ is observed only when $V=1$. As a result, proving the consistency of the estimator is nontrivial. Nevertheless, it can be ensured under the following identifiability assumption.
Assumption 5.2.
If $f_{1}(t\mid\bm{x};\bm{\theta}_{1})=f_{1}(t\mid\bm{x};\bm{\theta}_{1}^{\prime})$ and $f_{2}(t\mid\bm{x};\bm{\theta}_{2})=f_{2}(t\mid\bm{x};\bm{\theta}_{2}^{\prime})$ for any $t\in[0,\infty)$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$.
Theorem 5.3.
Under assumption 5.2, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ without administrative censoring is asymptotically consistent.
As in standard survival models, the distribution of the censoring time can be ignored when estimating the hazard function of the event time. Notably, section D.1 shows that this property remains valid even when the censoring times for the earlier and later events differ.
When the hazard function of $C$ is also of interest and $(\bm{\theta}_{1},\bm{\theta}_{2})$ shares some components with $\bm{\theta}_{c}$, the following assumption facilitates the joint estimation of $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$. However, it should be noted that this additional assumption holds only for restrictive survival models such as the constant hazard function.
Assumption 5.4.
If $\bm{\theta}_{c}\neq\bm{\theta}_{c}^{\prime}$, then the sign of $S_{c}(t\mid\bm{x};\bm{\theta}_{c})-S_{c}(t\mid\bm{x};\bm{\theta}_{c}^{\prime})$ is nonzero and constant over $t\in[0,\infty)$.
Proposition 5.5.
Under assumption 5.2 and assumption 5.4, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}(\bm{\theta}_{c})$ without administrative censoring is asymptotically consistent.
Next, we consider the case with administrative censoring. The structure of the probability model differs slightly from that described in lemma 5.1.
Lemma 5.6.
For $0\leq y\leq\tau$ and $v=0$, the joint distribution of $(\bar{Y},\bar{V})$ is
$\bar{g}_{v=0}(y)\equiv\begin{cases}S_{\circ}(\tau)S_{c}(\tau)&(y=\tau)\\ S_{\circ}(y)f_{c}(y)&(y\neq\tau).\end{cases}$ (9)
For $0\leq z\leq y\leq\tau$ and $v=1$, the joint distribution of $(\bar{Z},\bar{Y},\bar{V})$ is $g_{v=1}(z,y)$.
As in lemma 5.6, the presence of administrative censoring results in the replacement of the density $f_{c}$ in $g_{v=0}$ at $y=\tau$ by the survival function $S_{c}$. Then, the log likelihood is given by $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})$. This log likelihood appears to allow for the estimation of the hazard functions of $T_{1}$ and $T_{2}$ without requiring the censoring time distribution. However, since observations are only available up to time $t=\tau$, we cannot impose the condition that the integrals of $f_{1}$ and $f_{2}$ over $t\in[0,\infty)$ equal one. This condition is used in the proof of theorem 5.3 to ensure the identifiability of the survival model in the absence of administrative censoring. The following theorem shows that alternative assumptions can nevertheless guarantee the consistency of the estimator.
Assumption 5.7.
If $f_{1}(t\mid\bm{x};\bm{\theta}_{1})=f_{1}(t\mid\bm{x};\bm{\theta}_{1}^{\prime})$ for any $t\in[0,\tau]$ and $\bm{x}\in\mathcal{X}$, then $\bm{\theta}_{1}=\bm{\theta}_{1}^{\prime}$.
Assumption 5.8.
$\bm{\theta}_{2}$ is known and denoted by $\bm{\theta}_{2}^{*}$.
Theorem 5.9.
Under assumption 5.7 and assumption 5.8, the estimator of $\bm{\theta}_{1}$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}^{*})$ with administrative censoring is asymptotically consistent.
The following identifiability on the censoring time distribution allows for the joint estimation of $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$.
Assumption 5.10.
If $f_{1}(t\mid\bm{x};\bm{\theta}_{1})=f_{1}(t\mid\bm{x};\bm{\theta}_{1}^{\prime})$ and $f_{c}(t\mid\bm{x};\bm{\theta}_{c})=f_{c}(t\mid\bm{x};\bm{\theta}_{c}^{\prime})$ for any $t\in[0,\tau]$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{c})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{c}^{\prime})$.
Proposition 5.11.
Under assumption 5.8 and assumption 5.10, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}^{*})+\mathcal{L}_{c}(\bm{\theta}_{c})$ with administrative censoring is asymptotically consistent.
Unfortunately, theorem 5.9 and proposition 5.11 rely on assumption 5.8, which is unrealistic in many practical applications. Furthermore, assumption 5.7 and assumption 5.10 are also overly restrictive when the administrative censoring time $\tau$ is short. The two-stage estimation approach proposed in section 5.3 addresses this issue arising from administrative censoring. By transferring parameters estimated from a source domain that is not subject to administrative censoring, the method avoids reliance on these restrictive assumptions. Note that the estimation procedure for the case in which the administrative censoring time $\tau$ is sufficiently long is described in section D.2.
5.2 Parameter Estimation
In the absence of administrative censoring, one approach to jointly estimate the parameters of the earlier and later hazard functions is to directly maximize the objective function $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$. However, computing its gradient with respect to the parameters is often challenging, as $\log S_{\circ}(y)$ does not admit a closed-form expression. Consequently, instead of directly maximizing the objective function, we adopt the following lower bound derived using Jensen’s inequality (Billingsley 2012).
Theorem 5.12.
Define $\mathcal{L}_{1}(\bm{\theta}_{1}\mid\mathcal{D}_{1})$ as the log likelihood obtained by replacing $f$ and $\mathcal{D}$ in eq. 3 with $f_{1}$ and $\mathcal{D}_{1}\equiv(\bm{x}_{i},z_{i},w_{i})_{i=1}^{n}$, respectively. Similarly, define $\mathcal{L}_{2}(\bm{\theta}_{2}\mid\mathcal{D}_{2})$ by replacing $f$ and $\mathcal{D}$ in eq. 3 with $f_{2}$ and $\mathcal{D}_{2}\equiv(\bm{x}_{i},y_{i}-z_{i},v_{i})_{i=1}^{n}$, respectively. Then, $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ is bounded below by
$\mathcal{I}\bigl[\bigl(\mathcal{L}_{1}(\bm{\theta}_{1}\mid\mathcal{D}_{1})+\mathcal{L}_{2}(\bm{\theta}_{2}\mid\mathcal{D}_{2})-\log q\bigr)q\bigr],$ (10)
where $q\equiv\prod_{i=1}^{m}q_{i}(z_{i},w_{i})$, $q_{i}$ is a positive function such that $\mathcal{I}_{i}[q_{i}(z_{i},w_{i})]=1$, $\mathcal{I}\equiv\mathcal{I}_{1}\circ\cdots\circ\mathcal{I}_{m}$, and $\mathcal{I}_{i}$ is given by
$\mathcal{I}_{i}[~\cdot~]\equiv[~\cdot~]\big|_{z_{i}=y_{i},w_{i}=0}+\int_{0}^{y_{i}}[~\cdot~]\big|_{w_{i}=1}dz_{i}.$ (11)
Given the parameters $(\bm{\theta}_{1},\bm{\theta}_{2})$, the lower bound defined in eq. 10 is maximized when $q_{i}(z_{i},w_{i})$ is given by
$\begin{cases}S_{\circ}(y_{i})^{-1}S_{1}(z_{i}\mid\bm{x}_{i})&(w_{i}=0)\\ S_{\circ}(y_{i})^{-1}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})&(w_{i}=1).\end{cases}$ (12)
Based on theorem 5.12, the expectation-maximization algorithm (Dempster, Laird, and Rubin 1977) is implemented as in algorithm 1, iteratively updating $q$ and optimizing the parameters accordingly.
The integral in eq. 10 cannot be evaluated in closed form. To resolve this issue, we adopt a Monte Carlo approximation as proposed by Wei and Tanner (1990). The missing values in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are imputed using samples generated from $q$. This procedure is repeated $s$ times, and the resulting pseudo-complete datasets are then combined. Let $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ denote the aggregated pseudo-complete datasets corresponding to $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, respectively. The lower bound is approximated by
$s^{-1}\mathcal{L}_{1}(\bm{\theta}_{1}\mid\mathcal{P}_{1})+s^{-1}\mathcal{L}_{2}(\bm{\theta}_{2}\mid\mathcal{P}_{2})-\mathcal{I}[q\log q].$ (13)
The formulation in eq. 13 enables the independent optimization of $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$ using standard full-likelihood methods, thereby facilitating the reuse of existing resources and models. This approximation can also be applied to maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}^{*})$ in the presence of administrative censoring.
The remaining problem is how to efficiently generate samples from $q$ for use in Monte Carlo integration. As shown in eq. 12, the normalization constant $S_{\circ}(y_{i})$ appears in the sampling distribution $q_{i}(z_{i},w_{i})$. To avoid computing this constant explicitly, we adopt the rejection sampling scheme described in algorithm 2. The relative error of the gradient of eq. 10 with respect to the parameters is $\mathcal{O}\left(\frac{1}{n}\sqrt{\frac{m}{s}}\right)$.
Algorithm 1 Lower Bound Maximization
1:while iteration count has not reached the maximum do
2:  Update $q$ to the form defined in eq. 12.
3:  Generate pseudo-complete datasets using $q$.
4:  Optimize the parameters based on eq. 13.
5:end while
Algorithm 2 Reject Sampling
1: Repeat $t\sim f_{1}(t\mid\bm{x}_{i})$ until one of the following holds:
2:if the accident time $t$ exceeds the censoring time $y_{i}$ then
3:  return the imputed value $(z_{i},w_{i})\leftarrow(y_{i},\ 0)$
4:else if $\text{Uniform}(0,1)\leq S_{2}(y_{i}-t\mid\bm{x}_{i})$ then
5:  return the imputed value $(z_{i},w_{i})\leftarrow(t,\ 1)$
6:end if
5.3 Two-Stage Estimation
Based on the results obtained thus far, we propose the two-stage estimation approach. Assume that the true earlier hazard functions in both the source and target domains follow proportional hazards models, and that they share the same baseline hazard function and delay distribution. As described in section 4, the source domain is assumed to be free from administrative censoring, whereas the target domain is subject to it.
In the first stage, the earlier and later hazard functions are jointly estimated from the source domain data. The earlier hazard function in the source domain is specified as $h_{b}(t;\bm{\alpha})\phi(\bm{x};\bm{\beta})$ for $t\in[0,\infty)$ and $\bm{x}\in\mathcal{X}$, where both $h_{b}$ and $\phi$ are strictly positive functions parameterized by unknown real-valued vectors $\bm{\alpha}$ and $\bm{\beta}$, respectively. Let $(\hat{\bm{\alpha}},\hat{\bm{\beta}},\hat{\bm{\theta}}_{2})$ denote the estimator of $(\bm{\alpha},\bm{\beta},\bm{\theta}_{2})$ obtained by maximizing $\mathcal{E}_{s}((\bm{\alpha},\bm{\beta}),\bm{\theta}_{2})$, where $\mathcal{E}_{s}$ refers to $\mathcal{E}$ evaluated over the observations from the source domain. The joint parameter estimation can be achieved using algorithm 1.
[…]
Theorem 5.13.
Suppose that assumption 5.2 holds in the first stage. Then, $\hat{\gamma}$ is asymptotically consistent.
The expression of $\hat{\gamma}$ is not available in closed form. Although algorithm 1 is generally stable for estimating the single parameter $\gamma$ in a given cohort, repeatedly applying it across multiple cohorts can be computationally expensive. We therefore consider constructing a simple approximation of $\hat{\gamma}$. In the following, $(\mathcal{D}_{1},\mathcal{D}_{2})$ is regarded as the observations from the target domain. As a naive approach, we estimate $\gamma$ by ignoring the second term in $S_{\circ}(y)$.
Proposition 5.14.
Let $\check{\gamma}_{0}$ be defined as
$\frac{n-m}{\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t;\hat{\bm{\alpha}})dt},$ (14)
where $y_{i}^{\dagger}\equiv(1-v_{i})y_{i}+v_{i}z_{i}$. Then, $\check{\gamma}_{0}$ is less than $\hat{\gamma}$.
The second term in $S_{\circ}(y)$ captures the probability that an accident occurs but is not reported. Although a closed-form approximation of $\gamma$ can be obtained by omitting this term, proposition 5.14 shows that such an omission leads to an underestimation. To address this issue, we analyze $\hat{\gamma}$ by incorporating the difference between the survival probabilities of $T_{1}$ and $T_{2}$. For $\delta\in[0,y_{i}]$, let $\epsilon_{i}$ be defined as
$\epsilon_{i}(\delta)\equiv\int_{0}^{y_{i}-\delta}\frac{S_{2}(y_{i}-t\mid\bm{x}_{i};\hat{\bm{\theta}}_{2})}{\exp\bigl(-(y_{i}-t)\nu_{i}\bigr)}dt,$ (15)
[…]
Assumption 5.15.
$y_{i}\nu_{i}\ll 1$ for each $1\leq i\leq m$.
Theorem 5.16.
Let $\check{\gamma}$ be defined as
$\frac{n-m}{\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t;\hat{\bm{\alpha}})\bigl(1-S_{2}(y_{i}-t\mid\bm{x}_{i};\hat{\bm{\theta}}_{2})\bigr)^{1-v_{i}}dt}.$ (16)
Under assumption 5.15, the following holds:
$\frac{\hat{\gamma}}{\check{\gamma}}=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right),$ (17)
where $\delta_{i}^{*}$ is the fixed point satisfying $\delta_{i}^{*}=\epsilon_{i}(\delta_{i}^{*})$.
Intuitively, the denominator in eq. 16 can be regarded as a weighted version of that in eq. 14, where the weights correspond to the probability that an accident occurring at each time is reported within the remaining observation period. In many applications, the hazard of reporting delays is substantially greater than that of accident occurrences. In such cases, $\delta_{i}^{*}$ becomes sufficiently small relative to $y_{i}$ because $\epsilon_{i}(\delta)$ exhibits a steep decline. Consequently, under assumption 5.15, $\check{\gamma}$ provides a close approximation to $\hat{\gamma}$.
[…]
Table 1: Parameter estimation results on the toy model. The table reports the mean and standard deviation of estimated parameters across $100$ trials. Cells with a thin gray background indicate that the relative error of the mean from the true value exceeds 55%, while those with a darker gray background indicate an error exceeding 2525%. Let $\lambda_{*}$ denote the true value of $\lambda$.
6 Numerical Study
In this section, we provide experimental evidence supporting the validity of our approach.
Survival Model. The earlier hazard function follows a parametric proportional hazards model. Its baseline hazard function is defined as the piecewise constant function $h_{b}(t)=\sum_{i=1}^{3}\alpha_{i}\mathbb{I}\bigl(t\in[\xi_{i-1},\xi_{i})\bigr)$, where $(\xi_{0},\xi_{1},\xi_{2},\xi_{3})=(0,0.5,1,\infty)$ and each $\alpha_{i}\in(0,\infty)$. Note that the theoretical behavior of the proportional hazards model with this baseline is detailed in Friedman (1982). The covariate effect in the source domain is modeled as $\exp([\beta_{1},\cdots,\beta_{d}]^{\top}\bm{x})$ for $\bm{x}\in\mathcal{X}$, where each $\beta_{i}\in\mathbb{R}$. In contrast, the covariate effect in the target domain is represented by a scalar parameter $\gamma$. The later hazard function follows a constant hazard rate $\lambda$.
Estimation. The parameters for the source domain were estimated using algorithm 1 with $30$ iterations and $s=10$. The sampling of the missing values $(z_{i},w_{i})_{i=1}^{m}$ was performed via algorithm 2. The parameters were initialized using estimates obtained by substituting the missing values with $(y_{i},v_{i})_{i=1}^{m}$. For the target domain, $\gamma$ was estimated by eq. 16. We refer to this estimation procedure as Ours. One baseline for both the source and target domains assumes that the missing values are fully observed, while another replaces them with $(y_{i},v_{i})_{i=1}^{m}$. These baselines are referred to as Complete and Naive, respectively.
Environment. We used a 64-bit Windows machine with the Intel Core i9-9900K @ $3.60$ GHz processor and $64$ GB of RAM. The code was implemented by Python version 3.13.5. The maximization of $\mathcal{L}_{1}(\bm{\theta}_{1}\mid\mathcal{P}_{1})$ was conducted using version 0.30.0 of the lifelines library [^1] https://github.com/CamDavidsonPilon/lifelines. All parameters were set to their default values.
6.1 Toy Model
We investigate the behavior of our parameter estimation.
Setting. We generated $n=1000$ datasets for each of the source and target domains based on the survival model described above. The true value of $(\alpha_{1},\alpha_{2},\alpha_{3},\beta_{1},\gamma)$ was set to $(0.1,0.2,0.3,1,2)$. To assess the effect of reporting delays, we considered three distinct true values of $\lambda$: $0.5$, 55, and $50$. The censoring time followed a constant hazard rate of $1$. The administrative censoring time in the target domain was set to $\tau=0.75$. Each covariate $\bm{x}$ with $d=1$ for the sorce domain was generated uniformly over [1,1][-1,1]. To clarify the role of the second stage in the two-stage estimation procedure, we used the true values of $(\alpha_{1},\alpha_{2},\alpha_{3},\lambda)$ when estimating $\gamma$ in this subsection only.
Result. Table 1 highlights that adjusting for reporting delays helps mitigate the underestimation caused by ignoring them. Notably, the parameter estimates from our approach closely match the true values, except in cases with an excessive number of long delays. Additional results for finer values of $\lambda_{*}$ are provided in section E.1.
6.2 Insurance Risk Modeling
We demonstrate the practical utility of our approach for the timely evaluation of insurance risk.
Setting. To explore the potential of expanding the eligible age range for underwriting high-risk individuals, we adopted two benchmark datasets, Dialysis (Carvalho et al. 2003) and Support2 (Knaus et al. 1995), which are available via version 0.2.9 of the SurvSet library [^2] https://github.com/chl8856/SurvSet. These datasets comprise $6805$ and $9105$ individuals, respectively. The time and event indicators in these datasets were aligned to represent accident occurrences $(z_{i}^{\prime},w_{i}^{\prime})$, where $z_{i}^{\prime}$ was standardized by dividing each value by its median. In addition to the original censoring present in the datasets, synthetic censoring times $c_{i}$ were independently generated under a constant hazard rate of $1$. The observed accident time and indicator were then redefined as $z_{i}=\min\{z_{i}^{\prime},c_{i}\}$ and $w_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}\leq c_{i})$, respectively. Synthetic reporting times $r_{i}$ were simulated under a constant hazard rate of $\lambda_{*}=5$. Let $y_{i}=\min\{z_{i}^{\prime}+w_{i}r_{i},c_{i}\}$ and $v_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}+w_{i}r_{i}\leq c_{i})$. In the source domain, age was employed as the covariate. We denote by Young the setting in which the youngest 25% of the population is taken as the target domain, with the remaining 75% serving as the source domain. Conversely, Senior refers to the setting where the oldest 25% is taken as the target domain and the remaining 75% as the source domain. For each of the $100$ trials, the dataset from the target domain was randomly shuffled and evenly split into two subsets. One subset, administratively censored at $\tau=0.75$, was used for parameter estimation in the second stage of the two-stage procedure. The other subset, free from administrative censoring, was employed to evaluate the estimated survival model. This test set represents individuals expected to be underwritten as new policyholders at time $\tau$ after the target domain cohort has emerged. Each individual receives a unit benefit if an accident occurs, and the premium is calculated as the integral of the estimated hazard rate for accident occurrence over the observation window. The evaluation metric is defined as the total difference between paid benefits and collected premiums, normalized by the total premiums.
Result. Table 2 shows that accounting for reporting delays facilitates timely risk evaluation under administrative censoring, even before the complete dataset is available. The estimated parameters are presented in section E.2.
7 Conclusion
This study tackled learning survival models with right-censored reporting delays. Theoretical and experimental results highlight the effectiveness of our approach.
[] Dataset Target Method Metric Dialysis Young Complete $0.044\pm 0.180$ Dialysis Young Naive $0.362\pm 0.270$ Dialysis Young Ours $0.085\pm 0.181$ Dialysis Senior Complete $0.003\pm 0.112$ Dialysis Senior Naive $0.366\pm 0.173$ Dialysis Senior Ours $0.086\pm 0.113$ Support2 Young Complete $0.023\pm 0.077$ Support2 Young Naive $0.281\pm 0.109$ Support2 Young Ours $0.029\pm 0.082$ Support2 Senior Complete $0.009\pm 0.067$ Support2 Senior Naive $0.322\pm 0.096$ Support2 Senior Ours $0.010\pm 0.072$
Table 2: Performance in insurance risk modeling. The mean and standard deviation of the evaluation metric across $100$ trials are reported; closer to zero is better.
Appendix A Notations
[] Notation Description $\mathcal{D}$ $\mathcal{D}\equiv(\bm{x}_{i},y_{i},v_{i})_{i=1}^{n}$ is the set of independent observations $\mathcal{D}_{1}$ $\mathcal{D}_{1}\equiv(\bm{x}_{i},z_{i},w_{i})_{i=1}^{n}$ is the set of independent observations for the earlier process $\mathcal{D}_{2}$ $\mathcal{D}_{2}\equiv(\bm{x}_{i},y_{i}-z_{i},v_{i})_{i=1}^{n}$ is the set of independent observations for the later process $\mathcal{P}_{1}$ the combined pseudo-complete dataset corresponding for the earlier process $\mathcal{P}_{2}$ the combined pseudo-complete dataset corresponding for the later process $n$ the number of observations $m$ the number of observations with $v_{i}=0$ $s$ the number of pseudo-complete datasets $Y$ $\min\{T,C\}$ $V$ $\mathbb{I}(T\leq C)$ $\bar{Y}$ $\min\{Y,\tau\}$ $\bar{V}$ $\mathbb{I}(T\leq C,T\leq\tau)$ $Z$ $\min\{T_{1},C\}$ $W$ $\mathbb{I}(T_{1}\leq C)$ $\bar{Z}$ $\min\{Z,\tau\}$ $\bar{W}$ $\mathbb{I}(T_{1}\leq C,T_{1}\leq\tau)$ $\bm{x}_{i}$ $\bm{x}_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ represents the individual covariates $y_{i}$ $y_{i}\in[0,\infty)$ is the time from the start of the earlier process to the later event or censoring $v_{i}$ $v_{i}\in\{0,1\}$ is the later event indicator $z_{i}$ $z_{i}\in[0,y_{i}]$ is the time until the earlier event or censoring $w_{i}$ $w_{i}\in\{0,1\}$ is the earlier event indicator $\mathcal{L}(\bm{\theta})$ $\log\prod_{i=1}^{n}f(y_{i}\mid\bm{x}_{i})^{v_{i}}S(y_{i}\mid\bm{x}_{i})^{1-v_{i}}$ $\mathcal{L}_{c}(\bm{\theta}_{c})$ $\log\prod_{i=1}^{n}S_{c}(y_{i}\mid\bm{x}_{i})^{v_{i}}f_{c}(y_{i}\mid\bm{x}_{i})^{1-v_{i}}$ $\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})$ $\log\prod_{i=1}^{n}S_{c}(y_{i}\mid\bm{x}_{i})^{v_{i}^{\dagger}}f_{c}(y_{i}\mid\bm{x}_{i})^{1-v_{i}^{\dagger}}$ with $v_{i}^{\dagger}\equiv\mathbb{I}(y_{i}\geq\tau(1-v_{i}))$ $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ $\log\prod_{i=1}^{n}S_{\circ}(y_{i}\mid\bm{x}_{i})^{1-v_{i}}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})^{v_{i}}$ $\bm{\theta}$ the real-valued vector that parameterize the probability density function of $T$ $\bm{\theta}_{c}$ the real-valued vector that parameterize the probability density function of $C$ $\bm{\theta}_{1}$ the real-valued vector that parameterize the probability density function of $T_{1}$ $\bm{\theta}_{2}$ the real-valued vector that parameterize the probability density function of $T_{2}$ $f(t)$ the probability density function of $T$ at time $t$ $f_{c}(t)$ the probability density function of $C$ at time $t$ $f_{1}(t)$ the probability density function of $T_{1}$ at time $t$ $f_{2}(t)$ the probability density function of $T_{2}$ at time $t$ $S(t)$ $\int_{t}^{\infty}f(s)ds$ $S_{c}(t)$ $\int_{t}^{\infty}f_{c}(s)ds$ $S_{1}(t)$ $\int_{t}^{\infty}f_{1}(s)ds$ $S_{2}(t)$ $\int_{t}^{\infty}f_{2}(s)ds$ $f_{\circ}(z,y)$ $f_{1}(z)f_{2}(y-z)$ $S_{\circ}(y)$ $S_{1}(y)+\int_{0}^{y}f_{1}(t)S_{2}(y-t)dt$ $\mathcal{I}$ $\mathcal{I}_{1}\circ\cdots\circ\mathcal{I}_{m}$ with $\mathcal{I}_{i}\equiv[~\cdot~]\big|_{z_{i}=y_{i},w_{i}=0}+\int_{0}^{y_{i}}[~\cdot~]\big|_{w_{i}=1}dz_{i}$ $h(t)$ the hazard function of $T$ at time $t$ $h_{b}(t)$ the baseline hazard function in a proportional hazards model
Appendix B Preliminaries
B.1 Standard Survival Model without Administrative Censoring
For $y\in[0,\infty)$ and $v=0$, the following holds:
$\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)$ $=\frac{\partial}{\partial y}\Pr(C\leq y,C<T)$
$=\frac{\partial}{\partial y}\int_{0}^{y}dc\int_{c}^{\infty}dtf(t)f_{c}(c)$
$=\int_{y}^{\infty}f(t)f_{c}(y)dt$
$=S(y)f_{c}(y).$
For $y\in[0,\infty)$ and $v=1$, the following holds:
$\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)$ $=\frac{\partial}{\partial y}\Pr(T\leq y,T\leq C)$
$=\frac{\partial}{\partial y}\int_{0}^{y}dt\int_{t}^{\infty}dcf(t)f_{c}(c)$
$=\int_{y}^{\infty}f(y)f_{c}(c)dc$
$=f(y)S_{c}(y).$
Consequently, for $y\in[0,\infty)$ and $v\in\{0,1\}$, the joint distribution of $(Y,V)$ is
$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1).\end{cases}$
From the definition of $g$, we have
$\frac{\partial S_{c}(y)}{\partial y}=-\frac{g(y,0)}{S(y)},~~~\frac{\partial S(y)}{\partial y}=-\frac{g(y,1)}{S_{c}(y)}.$
Using the Cauchy-Lipschitz uniqueness theorem (Teschl 2012), this equation with $S(0)=S_{c}(0)=1$ has a unique solution. Therefore, if $(f,f_{c})$ is identifiable with respect to $(\bm{\theta},\bm{\theta}_{c})$, then $g$ is also identifiable with respect to $(\bm{\theta},\bm{\theta}_{c})$. According to the weak law of large numbers (Billingsley 2012) and Jensen’s inequality, the following holds:
$\frac{1}{n}\bigl(\mathcal{L}(\bm{\theta})+\mathcal{L}_{c}(\bm{\theta}_{c})\bigr)$ $=\frac{1}{n}\sum_{i=1}^{n}\log g(y_{i},v_{i}\mid\bm{x}_{i};\bm{\theta},\bm{\theta}_{c})$
$\xrightarrow{p}\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta},\bm{\theta}_{c})$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log\frac{g(y,v\mid\bm{x};\bm{\theta},\bm{\theta}_{c})}{g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})$
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*}),$
where $\bm{\theta}^{*}$ and $\bm{\theta}_{c}^{*}$ denote the true values of $\bm{\theta}$ and $\bm{\theta}_{c}$, respectively, and $\rho(\bm{x})$ denotes the true probability distribution of $X$. Similarly, the following holds:
$\frac{1}{n}\mathcal{L}(\bm{\theta})$ $=\frac{1}{n}\sum_{i=1}^{n}\log f(y_{i}\mid\bm{x}_{i};\bm{\theta})^{v_{i}}S(y_{i}\mid\bm{x}_{i};\bm{\theta})^{1-v_{i}}$
$\xrightarrow{p}\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta})^{v}S(y\mid\bm{x};\bm{\theta})^{1-v}$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log\frac{f(y\mid\bm{x};\bm{\theta})^{v}S(y\mid\bm{x};\bm{\theta})^{1-v}}{f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log\frac{g(y,v\mid\bm{x};\bm{\theta},\bm{\theta}_{c}^{*})}{g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}$
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}.$
Consider the equality condition in Jensen’s inequality. If $(f,f_{c})$ is identifiable with respect to $(\bm{\theta},\bm{\theta}_{c})$, then the estimator of $(\bm{\theta},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{L}(\bm{\theta})+\mathcal{L}_{c}(\bm{\theta}_{c})$ is asymptotically consistent. If $f$ is identifiable with respect to $\bm{\theta}$, then the estimator of $\bm{\theta}$ obtained by maximizing $\mathcal{L}(\bm{\theta})$ is asymptotically consistent.
B.2 Standard Survival Model with Administrative Censoring
For $y=\tau$ and $v=0$, the following holds:
$\Pr(\bar{Y}=y,\bar{V}=v)$ $=\Pr(\tau\leq C,\tau<T)$
$=\int_{\tau}^{\infty}dc\int_{\tau}^{\infty}dtf(t)f_{c}(c)$
$=S(\tau)S_{c}(\tau).$
For $y\in[0,\tau)$ and $v=0$, the following holds:
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S(y)f_{c}(y).$
For $y\in[0,\infty)$ and $v=1$, the following holds:
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=f(y)S_{c}(y).$
Consequently, for $y\in[0,\tau]$ and $v\in\{0,1\}$, the joint distribution of $(\bar{Y},\bar{V})$ is
$\bar{g}(y,v)\equiv\begin{cases}S(\tau)S_{c}(\tau)&(v=0,y=\tau)\\ S(y)f_{c}(y)&(v=0,y\neq\tau)\\ f(y)S_{c}(y)&(v=1).\end{cases}$
Consider the case where $g$ in section B.1 is replaced by $\bar{g}$. If $(f,f_{c})$ over the time interval $0,\tau]$ is identifiable with respect to $(\bm{\theta},\bm{\theta}_{c})$, then the estimator of $(\bm{\theta},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{L}(\bm{\theta})+\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})$ is asymptotically consistent. If $f$ over the time interval $0,\tau]$ is identifiable with respect to $\bm{\theta}$, the estimator of $\bm{\theta}$ obtained by maximizing $\mathcal{L}(\bm{\theta})$ is asymptotically consistent.
Appendix C Proofs
C.1 Proof of Lemma 5.1
For $0\leq y$ and $v=0$, the following holds:
$\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)$
$=\frac{\partial}{\partial y}\Pr(Y\leq y,W=0,V=0)+\frac{\partial}{\partial y}\Pr(Y\leq y,W=1,V=0)$
$=\frac{\partial}{\partial y}\Pr(C\leq y,C<T_{1})+\frac{\partial}{\partial y}\Pr(C\leq y,T_{1}\leq C,C<T_{1}+T_{2})$
$=\frac{\partial}{\partial y}\int_{0}^{y}dc\int_{c}^{\infty}dt_{1}f_{1}(t_{1})f_{c}(c)+\frac{\partial}{\partial y}\int_{0}^{y}dc\int_{0}^{c}dt_{1}\int_{c-t_{1}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c}(c)$
$=\int_{y}^{\infty}dt_{1}f_{1}(t_{1})f_{c}(y)+\int_{0}^{y}dt_{1}\int_{y-t_{1}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c}(y)$
$=S_{1}(y)f_{c}(y)+\int_{0}^{y}f_{1}(t_{1})S_{2}(y-t_{1})f_{c}(y)dt_{1}$
$=S_{\circ}(y)f_{c}(y).$
For $0\leq z\leq y$ and $v=1$, the following holds:
$\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)$
$=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,W=0,V=1)+\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,W=1,V=1)$
$=\frac{\partial^{2}}{\partial y\partial z}\Pr(C\leq z,T_{1}+T_{2}\leq y,C<T_{1},T_{1}+T_{2}\leq C)$
$~~~~+\frac{\partial^{2}}{\partial y\partial z}\Pr(T_{1}\leq z,T_{1}+T_{2}\leq y,T_{1}\leq C,T_{1}+T_{2}\leq C)$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}+t_{2}}^{\infty}dcf_{1}(t_{1})f_{2}(t_{2})f_{c}(c)$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dt_{1}\int_{0}^{y-t_{1}}dt_{2}f_{1}(t_{1})f_{2}(t_{2})S_{c}(t_{1}+t_{2})$
$=\frac{\partial}{\partial y}\int_{0}^{y-z}f_{1}(z)f_{2}(t_{2})S_{c}(z+t_{2})dt_{2}$
$=f_{1}(z)f_{2}(y-z)S_{c}(y)$
$=f_{\circ}(z,y)S_{c}(y).$
C.2 Proof of Theorem 5.3
The following holds:
$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$
Considering that both $f_{1}$ and $f_{2}$ are the probability density functions, we have
$f_{1}(0)=\int_{0}^{\infty}f_{\circ}(0,t)dt,~~~f_{2}(0)=\int_{0}^{\infty}f_{\circ}(t,t)dt.$
Therefore, using assumption 5.2, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$ for any $0\leq z\leq y$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$. From the weak law of large numbers and Jensen’s inequality, we have
$\frac{1}{n}\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ $=\frac{1}{n}\sum_{i=1}^{n}\bigl((1-v_{i})\log S_{\circ}(y_{i}\mid\bm{x}_{i};\bm{\theta}_{1},\bm{\theta}_{2})+v_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i};\bm{\theta}_{1},\bm{\theta}_{2})\bigr)$
$\xrightarrow{p}\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log S_{\circ}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{S_{\circ}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})}{S_{\circ}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})}{f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log S_{\circ}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{g_{v=0}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c}^{*})}{g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c}^{*})}{g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log S_{\circ}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log S_{\circ}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*}),$
where $\bm{\theta}_{1}^{*}$ and $\bm{\theta}_{2}^{*}$ denote the true values of $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, respectively. Consequently, from the equality condition in Jensen’s inequality, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ is asymptotically consistent.
C.3 Proof of Proposition 5.5
For $0\leq z\leq y$, the following holds:
$\frac{\partial}{\partial y}\frac{g_{v=0}(y)}{f_{c}(y)}$ $=\frac{\partial}{\partial y}\Bigl(S_{1}(y)+\int_{0}^{y}f_{1}(t)S_{2}(y-t)dt\Bigr)$
$=\frac{\partial}{\partial y}\Bigl(1-\int_{0}^{y}f_{1}(t)\bigl(1-S_{2}(y-t)\bigr)dt\Bigr)$
$=\frac{\partial}{\partial y}\Bigl(1-\int_{0}^{y}dtf_{1}(t)\int_{0}^{y-t}dsf_{2}(s)\Bigr)$
$=-\int_{0}^{y}f_{1}(t)f_{2}(y-t)dt$
$=-\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt.$
Therefore, we have
$\frac{\partial}{\partial y}\frac{g_{v=0}(y)}{f_{c}(y)}+\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt=0$
$\Rightarrow\frac{1}{f_{c}(y)}\frac{\partial g_{v=0}(y)}{\partial y}-\frac{g_{v=0}(y)}{f_{c}(y)^{2}}\frac{\partial f_{c}(y)}{\partial y}+\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt=0$
$\Rightarrow-\frac{\partial g_{v=0}(y)}{\partial y}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{-1}+g_{v=0}(y)\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{-2}\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}+\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt=0$
$\Rightarrow\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}=\frac{\partial\log g_{v=0}(y)}{\partial y}\frac{\partial S_{c}(y)}{\partial y}-\frac{\int_{0}^{y}g_{v=1}(t,y)dt}{g_{v=0}(y)}\frac{1}{S_{c}(y)}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{2}.$
From the Cauchy-Lipschitz uniqueness theorem, $S_{c}$ is uniquely determined by specifying $f_{c}(0)$ given $(g_{v=0},g_{v=1})$. Here, the following holds:
$f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{f_{1}(0)S_{c}(t)}.$
Since both $f_{1}$ and $f_{2}$ are the probability density functions, we have
$g_{v=1}(0,0)=f_{1}(0)f_{2}(0)S_{c}(0)=\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,t)}{S_{c}(t)}dt\Bigr)\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(t,t)}{S_{c}(t)}dt\Bigr).$
From assumption 5.4, the right side of this equation monotonically increasing with respect to $f_{c}(0)$. Therefore, $S_{c}$ is uniquely determined given $(g_{v=0},g_{v=1})$. Additionally, the following holds:
$f_{1}(t)=\frac{g_{v=1}(t,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,s)}{S_{c}(s)}ds\Bigr)^{-1},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(s,s)}{S_{c}(s)}ds\Bigr)^{-1}.$
Therefore, applying assumption 5.2 and assumption 5.4, if $g_{v=0}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$ and $g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$ for any $0\leq z\leq y$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$. By the weak law of large numbers and Jensen’s inequality, we have
$\frac{1}{n}\bigl(\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}_{c}(\bm{\theta}_{c})\bigr)$ $=\frac{1}{n}\sum_{i=1}^{n}\bigl((1-v_{i})\log g_{v=0}(y_{i}\mid\bm{x}_{i};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})+v_{i}\log g_{v=1}(z_{i},y_{i}\mid\bm{x}_{i};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})\bigr)$
$\xrightarrow{p}\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=0}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$
$=\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{g_{v=0}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})}{g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log\frac{g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})}{g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})}$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})$
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})$
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*}).$
Consequently, applying the equality condition in Jensen’s inequality, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}_{c}(\bm{\theta}_{c})$ is asymptotically consistent.
C.4 Proof of Lemma 5.6
For $y=\tau$ and $v=0$, the following holds:
$\Pr(\bar{Y}=y,\bar{V}=v)$ $=\Pr(\bar{Y}=\tau,\bar{W}=0,\bar{V}=0)+\Pr(\bar{Y}=\tau,\bar{W}=1,\bar{V}=0)$
$=\Pr(\tau\leq C,\tau<T_{1})+\Pr(\tau\leq C,T_{1}\leq\tau,\tau<T_{1}+T_{2})$
$=\int_{\tau}^{\infty}dc\int_{\tau}^{\infty}dt_{1}f_{1}(t_{1})f_{c}(c)+\int_{\tau}^{\infty}dc\int_{0}^{\tau}dt_{1}\int_{\tau-t_{1}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c}(c)$
$=S_{1}(\tau)S_{c}(\tau)+\int_{0}^{\tau}f_{1}(t_{1})S_{2}(\tau-t_{1})S_{c}(\tau)dt_{1}$
$=S_{\circ}(\tau)S_{c}(\tau).$
For $0\leq y<\tau$ and $v=0$, the following holds:
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S_{\circ}(y)f_{c}(y).$
For $0\leq z\leq y\leq\tau$ and $v=1$, the following holds:
$\frac{\partial^{2}}{\partial y\partial z}\Pr(\bar{Z}\leq z,\bar{Y}\leq y,\bar{V}=v)=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)=f_{\circ}(z,y)S_{c}(y).$
C.5 Proof of Theorem 5.9
For $t\in[0,\tau]$, the following holds:
$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)}.$
Therefore, from assumption 5.7 and assumption 5.8, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2}^{*})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{*})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $\bm{\theta}_{1}=\bm{\theta}_{1}^{\prime}$. Consequently, replacing $g_{v=0}$ in section C.2 with $\bar{g}_{v=0}$, the estimator of $\bm{\theta}_{1}$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}^{*})$ is asymptotically consistent.
C.6 Proof of Proposition 5.11
For $t\in[0,\tau]$, the following holds:
$S_{c}(t)=\frac{f_{2}(0)g_{v=1}(0,t)}{g_{v=1}(0,0)f_{2}(t)},~~~f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)}.$
Therefore, from assumption 5.8 and assumption 5.10, if $g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2}^{*},\bm{\theta}_{c})=g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{\prime})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{c})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{c}^{\prime})$. Consequently, replacing $g_{v=0}$ in section C.3 with $\bar{g}_{v=0}$, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}^{*})+\mathcal{L}_{c}^{\dagger}(\bm{\theta}_{c})$ is asymptotically consistent.
C.7 Proof of Theorem 5.12
The following holds:
$\log\mathcal{I}\bigl[\exp\bigl(\mathcal{L}(\bm{\theta}_{1}\mid\mathcal{D}_{1})+\mathcal{L}(\bm{\theta}_{2}\mid\mathcal{D}_{2})\bigr)\bigr]$
$=\log\mathcal{I}\Bigl[\prod_{i=1}^{n}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}f_{2}(y_{i}-z_{i}\mid\bm{x}_{i})^{v_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})^{1-v_{i}}\Bigr]$
$=\log\mathcal{I}\Bigl[\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})\Bigr]\prod_{i=m+1}^{n}f_{1}(z_{i}\mid\bm{x}_{i})f_{2}(y_{i}-z_{i}\mid\bm{x}_{i})$
$=\log\prod_{i=1}^{m}\Bigl(S_{1}(y_{i}\mid\bm{x}_{i})+\int_{0}^{y_{i}}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})dz_{i}\Bigr)\prod_{i=m+1}^{n}f_{1}(z_{i}\mid\bm{x}_{i})f_{2}(y_{i}-z_{i}\mid\bm{x}_{i})$
$=\log\prod_{i=1}^{m}S_{\circ}(y_{i}\mid\bm{x}_{i})\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})$
$=\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}).$
Therefore, applying Jensen’s inequality, we have
$\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ $=\log\mathcal{I}\Bigl[\frac{\exp\bigl(\mathcal{L}(\bm{\theta}_{1}\mid\mathcal{D}_{1})+\mathcal{L}(\bm{\theta}_{2}\mid\mathcal{D}_{2})\bigr)}{q}q\Bigr]$
$\geq\mathcal{I}\bigl[\bigl(\mathcal{L}(\bm{\theta}_{1}\mid\mathcal{D}_{1})+\mathcal{L}(\bm{\theta}_{2}\mid\mathcal{D}_{2})-\log q\bigr)q\bigr]$
$=\mathcal{I}\Bigl[q\log\frac{\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})}{q}\Bigr]+\log\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i}).$
Additionally, for $1\leq i\leq m$, the following holds:
$\mathcal{I}_{i}\bigl[f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})\bigr]$ $=\bigl[f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})\bigr]\big|_{z_{i}=y_{i},w_{i}=0}$
$~~~~+\int_{0}^{y_{i}}\bigl[f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})\bigr]\big|_{w_{i}=1}dz_{i}$
$=S_{1}(y_{i}\mid\bm{x}_{i})+\int_{0}^{y_{i}}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})dz_{i}$
$=S_{\circ}(y_{i}).$
Consequently, according to the equality condition in Jensen’s inequality, the lower bound defined in eq. 10 is maximized when $q_{i}(z_{i},w_{i})$ is given by
$\begin{cases}S_{\circ}(y_{i})^{-1}S_{1}(z_{i}\mid\bm{x}_{i})&(w_{i}=0)\\ S_{\circ}(y_{i})^{-1}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})&(w_{i}=1).\end{cases}$
C.8 Proof of Theorem 5.13
According to theorem 5.3, we have
$\mathcal{E}_{t}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})\xrightarrow{p}\mathcal{E}_{t}((\bm{\alpha}_{*},\gamma),\bm{\theta}_{2}^{*}).$
Assumption 5.7 and assumption 5.8 are satisfied when the earlier and later hazard functions are given by $h_{b}(t;\bm{\alpha}_{*})\gamma$ and $h_{2}(t\mid\bm{x};\bm{\theta}_{2}^{*})$, respectively. Consequently, applying theorem 5.9, $\hat{\gamma}$ is asymptotically consistent.
C.9 Proof of Proposition 5.14
The following holds:
$\frac{\partial}{\partial\gamma}\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$
$=\frac{\partial}{\partial\gamma}\Bigl(\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\log f_{1}(z_{i}\mid\bm{x}_{i})f_{2}(y_{i}-z_{i}\mid\bm{x}_{i})\Bigr)$
$=\frac{\partial}{\partial\gamma}\Bigl(\sum_{i=1}^{m}\log\Bigl(S_{1}(y_{i}\mid\bm{x}_{i})+\int_{0}^{y_{i}}f_{1}(t\mid\bm{x}_{i})S_{2}(y_{i}-t\mid\bm{x}_{i})dt\Bigr)+\sum_{i=m+1}^{n}\log f_{1}(z_{i}\mid\bm{x}_{i})\Bigr)$
$=\frac{\partial}{\partial\gamma}\Bigl(\sum_{i=1}^{m}\Bigl(\log S_{1}(y_{i}\mid\bm{x}_{i})+\log\Bigl(1+\int_{0}^{y_{i}}\frac{f_{1}(t\mid\bm{x}_{i})S_{2}(y_{i}-t\mid\bm{x}_{i})}{S_{1}(y_{i}\mid\bm{x}_{i})}dt\Bigr)\Bigr)+\sum_{i=m+1}^{n}\log f_{1}(z_{i}\mid\bm{x}_{i})\Bigr)$
$=\frac{\partial}{\partial\gamma}\sum_{i=1}^{m}\Bigl(-\int_{0}^{y_{i}}h_{b}(t)\gamma dt+\log\Bigl(1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\Bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\Bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt\Bigr)\Bigr)$
$~~~~+\frac{\partial}{\partial\gamma}\sum_{i=m+1}^{n}\Bigl(\log h_{b}(z_{i})\gamma-\int_{0}^{z_{i}}h_{b}(t)\gamma dt\Bigr)$
$=-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\frac{n-m}{\gamma}+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(\gamma^{-1}+\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
Since the last term is positive, $\frac{\partial}{\partial\gamma}\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$ remains positive for any $0<\gamma\leq\check{\gamma}_{0}$. Consequently, $\check{\gamma}_{0}$ is less than $\hat{\gamma}$.
C.10 Proof of Theorem 5.16
Let $\eta$ be defined as
$\eta(\gamma)\equiv\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt-\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
Then, for $\delta_{i}\in[0,y_{i}]$, the following holds:
$|\eta(\gamma)|$ $\leq\sum_{i=1}^{m}\biggl|\Bigl(\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt\Bigr)\Bigl(1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\Bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\Bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt\Bigr)$
$~~~~~~~~~~~~-\int_{0}^{y_{i}}\Bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\Bigr)h_{b}(t)\exp\Bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\Bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt\biggr|$
$=\sum_{i=1}^{m}\int_{0}^{y_{i}}\Bigl(\exp\Bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\Bigr)-1\Bigr)h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~+\sum_{i=1}^{m}\int_{0}^{y_{i}}\Bigl(\int_{t}^{y_{i}}h_{b}(s)ds+\int_{0}^{y_{i}}h_{b}(s)S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)h_{b}(t)\gamma\exp\Bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\Bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$\leq\sum_{i=1}^{m}\int_{0}^{y_{i}}\bigl(\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)-1\bigr)\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~+\sum_{i=1}^{m}\int_{0}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$\leq\sum_{i=1}^{m}\int_{0}^{y_{i}-\delta_{i}}\bigl(\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)-1\bigr)\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\bigl(\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)-1\bigr)\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~+\sum_{i=1}^{m}\int_{0}^{y_{i}-\delta_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt.$
Therefore, we have
[…]
$~~~~+\sum_{i=1}^{m}\Bigl(y_{i}+\int_{0}^{y_{i}}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\sum_{i=1}^{m}\Bigl(\delta_{i}+\int_{0}^{y_{i}}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}^{2}\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)$
$=\sum_{i=1}^{m}\Bigl(\bigl(\exp(y_{i}\nu_{i})-1\bigr)\nu_{i}\epsilon_{i}(\delta_{i})+\bigl(\exp(\delta_{i}\nu_{i})-1\bigr)\delta_{i}\nu_{i}+y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}\exp\bigl(\delta_{i}\nu_{i}\bigr)$
$~~~~~~~~~~~~~~+\Bigl(\int_{0}^{y_{i}-\delta_{i}}S_{2}(y_{i}-s\mid\bm{x}_{i})ds+\int_{y_{i}-\delta_{i}}^{y_{i}}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\bigl(\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\nu_{i}^{2}\delta_{i}\exp(\delta_{i}\nu_{i})\bigr)\Bigr)$
$\leq\sum_{i=1}^{m}\Bigl(\bigl(\exp(y_{i}\nu_{i})-1\bigr)\nu_{i}\epsilon_{i}(\delta_{i})+\bigl(\exp(\delta_{i}\nu_{i})-1\bigr)\delta_{i}\nu_{i}+y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}\exp\bigl(\delta_{i}\nu_{i}\bigr)$
$~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr).$
Additionally, applying assumption 5.15, the following holds:
$)\displaystyle\mathcal{O}(\hat{\gamma}\eta(\hat{\gamma}$)) $=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(\bigl(\exp(y_{i}\nu_{i})-1\bigr)\nu_{i}\epsilon_{i}(\delta_{i})+\bigl(\exp(\delta_{i}\nu_{i})-1\bigr)\delta_{i}\nu_{i}+y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}\exp\bigl(\delta_{i}\nu_{i}\bigr)$
$~~~~~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}+y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}\bigl(1+\delta_{i}\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigl(1+\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(2y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+2\delta_{i}^{2}\nu_{i}^{2}+\delta_{i}^{3}\nu_{i}^{3}+\bigl(\epsilon_{i}(\delta_{i})^{2}+2\delta_{i}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}+\delta_{i}^{2}\nu_{i}\epsilon_{i}(\delta_{i})+\delta_{i}^{3}\nu_{i}\bigr)\nu_{i}^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(2y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+2\delta_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{3}\epsilon_{i}(\delta_{i})+3\delta_{i}^{2}\nu_{i}^{2}+2\delta_{i}^{3}\nu_{i}^{3}+\nu_{i}^{2}\epsilon_{i}(\delta_{i})^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(4y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+y_{i}^{2}\nu_{i}^{3}\epsilon_{i}(\delta_{i})+3\delta_{i}^{2}\nu_{i}^{2}+\nu_{i}^{2}\epsilon_{i}(\delta_{i})^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}+\nu_{i}^{2}\epsilon_{i}(\delta_{i})^{2}\Bigr)\Bigr)$
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl((\delta_{i}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)^{2}\Bigr)\Bigr).$
Here, the gradient of $\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$ can be represented as
$\frac{\partial}{\partial\gamma}\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$
$=\frac{n-m}{\gamma}-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
$~~~~-\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}$
$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}+\frac{n-m}{\check{\gamma}}-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)\bigl(1-(1-v_{i})S_{2}(y_{i}-t\mid\bm{x}_{i})\bigr)dt-\eta(\gamma)$
$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}-\eta(\gamma).$
Considering that this gradient is zero at $\gamma=\hat{\gamma}$, we have
$\frac{\hat{\gamma}}{\check{\gamma}}$ $=1-\frac{\hat{\gamma}\eta(\hat{\gamma})}{n-m}$
$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\Bigl((\delta_{i}^{*}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i}^{*})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i}^{*})\nu_{i}\bigr)^{2}\Bigr)}{n-m}\right)$
$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}(\delta_{i}^{*}\nu_{i})(y_{i}\nu_{i}+2\delta_{i}^{*}\nu_{i})}{n-m}\right)$
$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right).$
Appendix D Alternative Problem Setting
D.1 Double Censoring Times
Let $\tilde{Z}\equiv\min\{T_{1},C_{1}\},\tilde{Y}\equiv\min\{T_{1}+T_{2},C_{1}+C_{2}\},\tilde{W}\equiv\mathbb{I}(T_{1}\leq C_{1})$, and $\tilde{V}\equiv\mathbb{I}(T_{1}+T_{2}\leq C_{1}+C_{2})$, where $C_{1}$ and $C_{2}$ are non-negative random variables. Assume that $T_{1}$, $T_{2}$, and $(C_{1},C_{2})$ are mutually independent. Define $f_{c_{1},c_{2}}$ as the joint distribution of $(C_{1},C_{2})$.
Proposition D.1.
[…]
$S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$
[…]
$\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$
[…]
$f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
Proof.
For $0\leq y$ and $v=0$, the following holds:
$\frac{\partial}{\partial y}\Pr(\tilde{Y}\leq y,\tilde{V}=v)$
$=\frac{\partial}{\partial y}\Pr(C_{1}+C_{2}\leq y,C_{1}+C_{2}<T_{1}+T_{2})$
$=\frac{\partial}{\partial y}\int_{0}^{\infty}dc_{1}\int_{0}^{\max\{y-c_{1},0\}}dc_{2}\int_{0}^{\infty}dt_{1}\int_{\max\{c_{1}+c_{2}-t_{1},0\}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\frac{\partial}{\partial y}\int_{0}^{y}dc_{1}\int_{0}^{y-c_{1}}dc_{2}\int_{0}^{c_{1}+c_{2}}dt_{1}\int_{c_{1}+c_{2}-t_{1}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$~~~~+\frac{\partial}{\partial y}\int_{0}^{y}dc_{1}\int_{0}^{y-c_{1}}dc_{2}\int_{c_{1}+c_{2}}^{\infty}dt_{1}\int_{0}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\int_{0}^{y}dc_{1}\int_{0}^{y}dt_{1}\int_{y-t_{1}}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},y-c_{1})$
$~~~~+\int_{0}^{y}dc_{1}\int_{y}^{\infty}dt_{1}\int_{0}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},y-c_{1})$
$=S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$
For $0\leq z\leq y$ and $(w,v)=(0,1)$, the following holds:
$\frac{\partial^{2}}{\partial y\partial z}\Pr(\tilde{Z}\leq z,\tilde{Y}\leq y,\tilde{W}=w,\tilde{V}=v)$
$=\frac{\partial^{2}}{\partial y\partial z}\Pr(C_{1}\leq z,T_{1}+T_{2}\leq y,C_{1}<T_{1},T_{1}+T_{2}\leq C_{1}+C_{2})$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dc_{1}\int_{c_{1}}^{\infty}dt_{1}\int_{0}^{\max\{y-t_{1},0\}}dt_{2}\int_{t_{1}+t_{2}-c_{1}}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dc_{1}\int_{c_{1}}^{y}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}+t_{2}-c_{1}}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\frac{\partial}{\partial y}\int_{z}^{y}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}+t_{2}-z}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(z,c_{2})$
$=\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$
For $0\leq z\leq y$ and $(w,v)=(1,1)$, the following holds:
$\frac{\partial^{2}}{\partial y\partial z}\Pr(\tilde{Z}\leq z,\tilde{Y}\leq y,\tilde{W}=w,\tilde{V}=v)$
$=\frac{\partial^{2}}{\partial y\partial z}\Pr(T_{1}\leq z,T_{1}+T_{2}\leq y,T_{1}\leq C_{1},T_{1}+T_{2}\leq C_{1}+C_{2})$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}}^{\infty}dc_{1}\int_{\max\{t_{1}+t_{2}-c_{1},0\}}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}}^{t_{1}+t_{2}}dc_{1}\int_{t_{1}+t_{2}-c_{1}}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$~~~~+\frac{\partial^{2}}{\partial y\partial z}\int_{0}^{z}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}+t_{2}}^{\infty}dc_{1}\int_{0}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\frac{\partial}{\partial y}\int_{0}^{y-z}dt_{2}\int_{z}^{z+t_{2}}dc_{1}\int_{z+t_{2}-c_{1}}^{\infty}dc_{2}f_{1}(z)f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$~~~~+\frac{\partial}{\partial y}\int_{0}^{y-z}dt_{2}\int_{z+t_{2}}^{\infty}dc_{1}\int_{0}^{\infty}dc_{2}f_{1}(z)f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},c_{2})$
$=\int_{z}^{y}dc_{1}\int_{y-c_{1}}^{\infty}dc_{2}f_{1}(z)f_{2}(y-z)f_{c_{1},c_{2}}(c_{1},c_{2})+\int_{y}^{\infty}dc_{1}\int_{0}^{\infty}dc_{2}f_{1}(z)f_{2}(y-z)f_{c_{1},c_{2}}(c_{1},c_{2})$
$=f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
Proposition D.2.
Let $\tilde{\mathcal{E}}$ be defined as
$\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\Bigl(w_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})+(1-w_{i})\log\int_{z_{i}}^{y_{i}}f_{\circ}(t,y_{i}\mid\bm{x}_{i})dt\Bigr).$
The estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})$ is asymptotically consistent.
Proof.
This is evident in light of the discussion in section C.2. ∎
In this setting, the occurrence of an earlier event after the first censoring time is reported, but its timing is not observed. When $\Pr(C_{2}=0)=1$, this setting is equivalent to that described in section 4.
D.2 Assumption for Administrative Censoring
Since the survival probability of $T_{2}$ tends to be small for large values of the realization time, the following assumption appears reasonable when the administrative censoring time $\tau$ is sufficiently long.
Assumption D.3.
$S_{2}(\tau\mid\bm{x})=0$ for any $\bm{x}\in\mathcal{X}$.
Based on this assumption, we present the following proposition.
Proposition D.4.
Under assumption 5.7 and assumption D.3, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ with administrative censoring is asymptotically consistent.
Proof.
For $0\leq t\leq\tau$, the following holds:
$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$
From assumption D.3, we have
$f_{1}(0)=\int_{0}^{\tau}f_{\circ}(0,t)dt,~~~f_{2}(0)=\frac{f_{\circ}(0,0)}{f_{1}(0)}=\frac{f_{\circ}(0,0)}{\int_{0}^{\tau}f_{\circ}(0,t)dt}.$
Therefore, using assumption 5.7, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$. Based on the discussion in section C.2, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$ is asymptotically consistent. ∎
Appendix E Supplement to Numerical Study
E.1 Toy Model
[…]
Table 3: Supplementary results for the toy model. The display format is the same as in table 1.
E.2 Insurance Risk Modeling
[] Dataset Target Method Estimation Result $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\beta_{1}$ $\gamma$ Dialysis Young Complete $038\pm 0.004 0.0$$20\pm 0.003 0.0$$17\pm 0.002 0.0$$036\pm 0.002$ $4.044\pm 0.686$ Dialysis Young Naive $033\pm 0.005 0.0$$17\pm 0.003 0.0$$15\pm 0.003 0.0$$036\pm 0.002$ $3.298\pm 0.632$ Dialysis Young Ours $041\pm 0.006 0.0$$21\pm 0.003 0.0$$18\pm 0.003 0.0$$037\pm 0.002$ $4.006\pm 0.777$ Dialysis Senior Complete $061\pm 0.007 0.0$$34\pm 0.004 0.0$$24\pm 0.004 0.0$$028\pm 0.002$ $9.198\pm 1.204$ Dialysis Senior Naive $051\pm 0.008 0.0$$28\pm 0.005 0.0$$20\pm 0.004 0.0$$028\pm 0.003$ $8.147\pm 1.435$ Dialysis Senior Ours $067\pm 0.011 0.0$$37\pm 0.006 0.0$$27\pm 0.006 0.0$$028\pm 0.003$ $9.224\pm 1.668$ Support2 Young Complete $0.618\pm 0.041$ $149\pm 0.011 0.0$$87\pm 0.008 0.0$$012\pm 0.001$ $1.711\pm 0.144$ Support2 Young Naive $0.521\pm 0.050$ $128\pm 0.014 0.0$$75\pm 0.009 0.0$$011\pm 0.001$ $1.551\pm 0.177$ Support2 Young Ours $0.627\pm 0.062$ $152\pm 0.017 0.0$$90\pm 0.012 0.0$$012\pm 0.001$ $1.704\pm 0.201$ Support2 Senior Complete $0.659\pm 0.026$ $162\pm 0.009 0.0$$87\pm 0.008 0.0$$011\pm 0.001$ $2.529\pm 0.130$ Support2 Senior Naive $0.540\pm 0.031$ $134\pm 0.010 0.0$$73\pm 0.007 0.0$$011\pm 0.001$ $2.314\pm 0.173$ Support2 Senior Ours $0.663\pm 0.040$ $163\pm 0.012 0.0$$89\pm 0.009 0.0$$011\pm 0.001$ $2.473\pm 0.189$
Table 4: Supplementary results for insurance risk modeling. The table reports the mean and standard deviation of estimated parameters across $100$ trials. The estimated values of $\lambda$ were $4.532\pm 0.205$, $4.568\pm 0.270$, $4.914\pm 0.118$, and $4.933\pm 0.115$, in top-to-bottom order of the dataset and target shown in the table.

### v2

@@
 Note that the estimation procedure for the case in which the administrative censoring time $\tau$ is sufficiently long is described in section D.2.
-5.2 Parameter Estimation
+5.2 Parameter Optimization
 In the absence of administrative censoring, one approach to jointly estimate the parameters of the earlier and later hazard functions is to directly maximize the objective function $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})$.
@@
 2003) and Support2 (Knaus et al.
-1995), which are available via version 0.2.9 of the SurvSet library [^2] https://github.com/chl8856/SurvSet.
+1995), which are available via version 0.2.9 of the SurvSet library [^2] https://github.com/ErikinBC/SurvSet.
 These datasets comprise $6805$ and $9105$ individuals, respectively.
@@
 $=\int_{y}^{\infty}f(t)f_{c}(y)dt$
-$=S(y)f_{c}(y).$
+$=S(y)f_{c}(y).$ (18)
 For $y\in[0,\infty)$ and $v=1$, the following holds:
@@
 $=\int_{y}^{\infty}f(y)f_{c}(c)dc$
-$=f(y)S_{c}(y).$
+$=f(y)S_{c}(y).$ (19)
 Consequently, for $y\in[0,\infty)$ and $v\in\{0,1\}$, the joint distribution of $(Y,V)$ is
-$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1).\end{cases}$
+$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1).\end{cases}$ (20)
 From the definition of $g$, we have
-$\frac{\partial S_{c}(y)}{\partial y}=-\frac{g(y,0)}{S(y)},~~~\frac{\partial S(y)}{\partial y}=-\frac{g(y,1)}{S_{c}(y)}.$
+$\frac{\partial S_{c}(y)}{\partial y}=-\frac{g(y,0)}{S(y)},~~~\frac{\partial S(y)}{\partial y}=-\frac{g(y,1)}{S_{c}(y)}.$ (21)
 Using the Cauchy-Lipschitz uniqueness theorem (Teschl 2012), this equation with $S(0)=S_{c}(0)=1$ has a unique solution.
@@
 $~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})$
-$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*}),$
+$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*}),$ (22)
 where $\bm{\theta}^{*}$ and $\bm{\theta}_{c}^{*}$ denote the true values of $\bm{\theta}$ and $\bm{\theta}_{c}$, respectively, and $\rho(\bm{x})$ denotes the true probability distribution of $X$.
@@
 $~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}$
-$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}.$
+$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}.$ (23)
 Consider the equality condition in Jensen’s inequality.
@@
 $=\int_{\tau}^{\infty}dc\int_{\tau}^{\infty}dtf(t)f_{c}(c)$
-$=S(\tau)S_{c}(\tau).$
+$=S(\tau)S_{c}(\tau).$ (24)
 For $y\in[0,\tau)$ and $v=0$, the following holds:
-$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S(y)f_{c}(y).$
+$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S(y)f_{c}(y).$ (25)
 For $y\in[0,\infty)$ and $v=1$, the following holds:
-$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=f(y)S_{c}(y).$
+$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=f(y)S_{c}(y).$ (26)
 Consequently, for $y\in[0,\tau]$ and $v\in\{0,1\}$, the joint distribution of $(\bar{Y},\bar{V})$ is
-$\bar{g}(y,v)\equiv\begin{cases}S(\tau)S_{c}(\tau)&(v=0,y=\tau)\\ S(y)f_{c}(y)&(v=0,y\neq\tau)\\ f(y)S_{c}(y)&(v=1).\end{cases}$
+$\bar{g}(y,v)\equiv\begin{cases}S(\tau)S_{c}(\tau)&(v=0,y=\tau)\\ S(y)f_{c}(y)&(v=0,y\neq\tau)\\ f(y)S_{c}(y)&(v=1).\end{cases}$ (27)
 Consider the case where $g$ in section B.1 is replaced by $\bar{g}$.
@@
 $=S_{1}(y)f_{c}(y)+\int_{0}^{y}f_{1}(t_{1})S_{2}(y-t_{1})f_{c}(y)dt_{1}$
-$=S_{\circ}(y)f_{c}(y).$
+$=S_{\circ}(y)f_{c}(y).$ (28)
 For $0\leq z\leq y$ and $v=1$, the following holds:
@@
 $=f_{1}(z)f_{2}(y-z)S_{c}(y)$
-$=f_{\circ}(z,y)S_{c}(y).$
+$=f_{\circ}(z,y)S_{c}(y).$ (29)
 C.2 Proof of Theorem 5.3
 The following holds:
-$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$
+$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$ (30)
 Considering that both $f_{1}$ and $f_{2}$ are the probability density functions, we have
-$f_{1}(0)=\int_{0}^{\infty}f_{\circ}(0,t)dt,~~~f_{2}(0)=\int_{0}^{\infty}f_{\circ}(t,t)dt.$
+$f_{1}(0)=\int_{0}^{\infty}f_{\circ}(0,t)dt,~~~f_{2}(0)=\int_{0}^{\infty}f_{\circ}(t,t)dt.$ (31)
 Therefore, using assumption 5.2, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$ for any $0\leq z\leq y$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$.
@@
 $\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log S_{\circ}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*})$
-$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*}),$
+$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*}),$ (32)
 where $\bm{\theta}_{1}^{*}$ and $\bm{\theta}_{2}^{*}$ denote the true values of $\bm{\theta}_{1}$ and $\bm{\theta}_{2}$, respectively.
@@
 $=-\int_{0}^{y}f_{1}(t)f_{2}(y-t)dt$
-$=-\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt.$
+$=-\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt.$ (33)
 Therefore, we have
@@
 $\Rightarrow-\frac{\partial g_{v=0}(y)}{\partial y}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{-1}+g_{v=0}(y)\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{-2}\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}+\frac{1}{S_{c}(y)}\int_{0}^{y}g_{v=1}(t,y)dt=0$
-$\Rightarrow\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}=\frac{\partial\log g_{v=0}(y)}{\partial y}\frac{\partial S_{c}(y)}{\partial y}-\frac{\int_{0}^{y}g_{v=1}(t,y)dt}{g_{v=0}(y)}\frac{1}{S_{c}(y)}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{2}.$
+$\Rightarrow\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}=\frac{\partial\log g_{v=0}(y)}{\partial y}\frac{\partial S_{c}(y)}{\partial y}-\frac{\int_{0}^{y}g_{v=1}(t,y)dt}{g_{v=0}(y)}\frac{1}{S_{c}(y)}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{2}.$ (34)
 From the Cauchy-Lipschitz uniqueness theorem, $S_{c}$ is uniquely determined by specifying $f_{c}(0)$ given $(g_{v=0},g_{v=1})$. Here, the following holds:
-$f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{f_{1}(0)S_{c}(t)}.$
+$f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{f_{1}(0)S_{c}(t)}.$ (35)
 Since both $f_{1}$ and $f_{2}$ are the probability density functions, we have
-$g_{v=1}(0,0)=f_{1}(0)f_{2}(0)S_{c}(0)=\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,t)}{S_{c}(t)}dt\Bigr)\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(t,t)}{S_{c}(t)}dt\Bigr).$
+$g_{v=1}(0,0)=f_{1}(0)f_{2}(0)S_{c}(0)=\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,t)}{S_{c}(t)}dt\Bigr)\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(t,t)}{S_{c}(t)}dt\Bigr).$ (36)
 From assumption 5.4, the right side of this equation monotonically increasing with respect to $f_{c}(0)$.
@@
 Additionally, the following holds:
-$f_{1}(t)=\frac{g_{v=1}(t,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,s)}{S_{c}(s)}ds\Bigr)^{-1},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(s,s)}{S_{c}(s)}ds\Bigr)^{-1}.$
+$f_{1}(t)=\frac{g_{v=1}(t,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,s)}{S_{c}(s)}ds\Bigr)^{-1},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(s,s)}{S_{c}(s)}ds\Bigr)^{-1}.$ (37)
 Therefore, applying assumption 5.2 and assumption 5.4, if $g_{v=0}(y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$ and $g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$ for any $0\leq z\leq y$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime},\bm{\theta}_{c}^{\prime})$.
@@
 $\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\rho(\bm{x})g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=0}(y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})$
-$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*}).$
+$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*}).$ (38)
 Consequently, applying the equality condition in Jensen’s inequality, the estimator of $(\bm{\theta}_{1},\bm{\theta}_{2},\bm{\theta}_{c})$ obtained by maximizing $\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2})+\mathcal{L}_{c}(\bm{\theta}_{c})$ is asymptotically consistent.
@@
 $=S_{1}(\tau)S_{c}(\tau)+\int_{0}^{\tau}f_{1}(t_{1})S_{2}(\tau-t_{1})S_{c}(\tau)dt_{1}$
-$=S_{\circ}(\tau)S_{c}(\tau).$
+$=S_{\circ}(\tau)S_{c}(\tau).$ (39)
 For $0\leq y<\tau$ and $v=0$, the following holds:
-$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S_{\circ}(y)f_{c}(y).$
+$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S_{\circ}(y)f_{c}(y).$ (40)
 For $0\leq z\leq y\leq\tau$ and $v=1$, the following holds:
-$\frac{\partial^{2}}{\partial y\partial z}\Pr(\bar{Z}\leq z,\bar{Y}\leq y,\bar{V}=v)=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)=f_{\circ}(z,y)S_{c}(y).$
+$\frac{\partial^{2}}{\partial y\partial z}\Pr(\bar{Z}\leq z,\bar{Y}\leq y,\bar{V}=v)=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)=f_{\circ}(z,y)S_{c}(y).$ (41)
 C.5 Proof of Theorem 5.9
 For $t\in[0,\tau]$, the following holds:
-$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)}.$
+$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)}.$ (42)
 Therefore, from assumption 5.7 and assumption 5.8, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2}^{*})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{*})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $\bm{\theta}_{1}=\bm{\theta}_{1}^{\prime}$.
@@
 For $t\in[0,\tau]$, the following holds:
-$S_{c}(t)=\frac{f_{2}(0)g_{v=1}(0,t)}{g_{v=1}(0,0)f_{2}(t)},~~~f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)}.$
+$S_{c}(t)=\frac{f_{2}(0)g_{v=1}(0,t)}{g_{v=1}(0,0)f_{2}(t)},~~~f_{1}(t)=\frac{g_{v=1}(t,t)}{f_{2}(0)S_{c}(t)}.$ (43)
 Therefore, from assumption 5.8 and assumption 5.10, if $g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2}^{*},\bm{\theta}_{c})=g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{\prime})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{c})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{c}^{\prime})$.
@@
 $=\log\prod_{i=1}^{m}S_{\circ}(y_{i}\mid\bm{x}_{i})\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})$
-$=\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}).$
+$=\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}).$ (44)
 Therefore, applying Jensen’s inequality, we have
@@
 $\geq\mathcal{I}\bigl[\bigl(\mathcal{L}(\bm{\theta}_{1}\mid\mathcal{D}_{1})+\mathcal{L}(\bm{\theta}_{2}\mid\mathcal{D}_{2})-\log q\bigr)q\bigr]$
-$=\mathcal{I}\Bigl[q\log\frac{\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})}{q}\Bigr]+\log\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i}).$
+$=\mathcal{I}\Bigl[q\log\frac{\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})}{q}\Bigr]+\log\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i}).$ (45)
 Additionally, for $1\leq i\leq m$, the following holds:
@@
 $=S_{1}(y_{i}\mid\bm{x}_{i})+\int_{0}^{y_{i}}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})dz_{i}$
-$=S_{\circ}(y_{i}).$
+$=S_{\circ}(y_{i}).$ (46)
 Consequently, according to the equality condition in Jensen’s inequality, the lower bound defined in eq. 10 is maximized when $q_{i}(z_{i},w_{i})$ is given by
-$\begin{cases}S_{\circ}(y_{i})^{-1}S_{1}(z_{i}\mid\bm{x}_{i})&(w_{i}=0)\\ S_{\circ}(y_{i})^{-1}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})&(w_{i}=1).\end{cases}$
+$\begin{cases}S_{\circ}(y_{i})^{-1}S_{1}(z_{i}\mid\bm{x}_{i})&(w_{i}=0)\\ S_{\circ}(y_{i})^{-1}f_{1}(z_{i}\mid\bm{x}_{i})S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})&(w_{i}=1).\end{cases}$ (47)
 C.8 Proof of Theorem 5.13
 According to theorem 5.3, we have
-$\mathcal{E}_{t}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})\xrightarrow{p}\mathcal{E}_{t}((\bm{\alpha}_{*},\gamma),\bm{\theta}_{2}^{*}).$
+$\mathcal{E}_{t}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})\xrightarrow{p}\mathcal{E}_{t}((\bm{\alpha}_{*},\gamma),\bm{\theta}_{2}^{*}).$ (48)
 Assumption 5.7 and assumption 5.8 are satisfied when the earlier and later hazard functions are given by $h_{b}(t;\bm{\alpha}_{*})\gamma$ and $h_{2}(t\mid\bm{x};\bm{\theta}_{2}^{*})$, respectively.
@@
 $~~~~+\frac{\partial}{\partial\gamma}\sum_{i=m+1}^{n}\Bigl(\log h_{b}(z_{i})\gamma-\int_{0}^{z_{i}}h_{b}(t)\gamma dt\Bigr)$
-$=-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\frac{n-m}{\gamma}+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(\gamma^{-1}+\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
+$=-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\frac{n-m}{\gamma}+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(\gamma^{-1}+\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$ (49)
 Since the last term is positive, $\frac{\partial}{\partial\gamma}\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$ remains positive for any $0<\gamma\leq\check{\gamma}_{0}$.
@@
 Let $\eta$ be defined as
-$\eta(\gamma)\equiv\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt-\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
+$\eta(\gamma)\equiv\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt-\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$ (50)
 Then, for $\delta_{i}\in[0,y_{i}]$, the following holds:
@@
 $~~~~+\sum_{i=1}^{m}\int_{0}^{y_{i}-\delta_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt$
-$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt.$
+$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt.$ (51)
 Therefore, we have
@@
 $\leq\sum_{i=1}^{m}\Bigl(\bigl(\exp(y_{i}\nu_{i})-1\bigr)\nu_{i}\epsilon_{i}(\delta_{i})+\bigl(\exp(\delta_{i}\nu_{i})-1\bigr)\delta_{i}\nu_{i}+y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}\exp\bigl(\delta_{i}\nu_{i}\bigr)$
-$~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr).$
+$~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr).$ (52)
 Additionally, applying assumption 5.15, the following holds:
@@
 $=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl(y_{i}\nu_{i}^{2}\epsilon_{i}(\delta_{i})+\delta_{i}^{2}\nu_{i}^{2}+\nu_{i}^{2}\epsilon_{i}(\delta_{i})^{2}\Bigr)\Bigr)$
-$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl((\delta_{i}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)^{2}\Bigr)\Bigr).$
+$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl((\delta_{i}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)^{2}\Bigr)\Bigr).$ (53)
 Here, the gradient of $\mathcal{E}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})$ can be represented as
@@
 $=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}+\frac{n-m}{\check{\gamma}}-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)\bigl(1-(1-v_{i})S_{2}(y_{i}-t\mid\bm{x}_{i})\bigr)dt-\eta(\gamma)$
-$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}-\eta(\gamma).$
+$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}-\eta(\gamma).$ (54)
 Considering that this gradient is zero at $\gamma=\hat{\gamma}$, we have
@@
 $=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}(\delta_{i}^{*}\nu_{i})(y_{i}\nu_{i}+2\delta_{i}^{*}\nu_{i})}{n-m}\right)$
-$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right).$
+$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right).$ (55)
 Appendix D Alternative Problem Setting
@@
 […]
-$S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$
+$S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$ (56)
 […]
-$\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$
+$\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$ (57)
 […]
-$f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
+$f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$ (58)
 Proof.
@@
 $~~~~+\int_{0}^{y}dc_{1}\int_{y}^{\infty}dt_{1}\int_{0}^{\infty}dt_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(c_{1},y-c_{1})$
-$=S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$
+$=S_{\circ}(y)\int_{0}^{y}f_{c_{1},c_{2}}(c_{1},y-c_{1})dc_{1}.$ (59)
 For $0\leq z\leq y$ and $(w,v)=(0,1)$, the following holds:
@@
 $=\frac{\partial}{\partial y}\int_{z}^{y}dt_{1}\int_{0}^{y-t_{1}}dt_{2}\int_{t_{1}+t_{2}-z}^{\infty}dc_{2}f_{1}(t_{1})f_{2}(t_{2})f_{c_{1},c_{2}}(z,c_{2})$
-$=\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$
+$=\int_{z}^{y}f_{\circ}(t,y)dt\int_{y-z}^{\infty}f_{c_{1},c_{2}}(z,c_{2})dc_{2}.$ (60)
 For $0\leq z\leq y$ and $(w,v)=(1,1)$, the following holds:
@@
 $=\int_{z}^{y}dc_{1}\int_{y-c_{1}}^{\infty}dc_{2}f_{1}(z)f_{2}(y-z)f_{c_{1},c_{2}}(c_{1},c_{2})+\int_{y}^{\infty}dc_{1}\int_{0}^{\infty}dc_{2}f_{1}(z)f_{2}(y-z)f_{c_{1},c_{2}}(c_{1},c_{2})$
-$=f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
+$=f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$ (61)
 Proposition D.2.
 Let $\tilde{\mathcal{E}}$ be defined as
-$\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\Bigl(w_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})+(1-w_{i})\log\int_{z_{i}}^{y_{i}}f_{\circ}(t,y_{i}\mid\bm{x}_{i})dt\Bigr).$
+$\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\Bigl(w_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})+(1-w_{i})\log\int_{z_{i}}^{y_{i}}f_{\circ}(t,y_{i}\mid\bm{x}_{i})dt\Bigr).$ (62)
 The estimator of $(\bm{\theta}_{1},\bm{\theta}_{2})$ obtained by maximizing $\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})$ is asymptotically consistent.
@@
 For $0\leq t\leq\tau$, the following holds:
-$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$
+$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$ (63)
 From assumption D.3, we have
-$f_{1}(0)=\int_{0}^{\tau}f_{\circ}(0,t)dt,~~~f_{2}(0)=\frac{f_{\circ}(0,0)}{f_{1}(0)}=\frac{f_{\circ}(0,0)}{\int_{0}^{\tau}f_{\circ}(0,t)dt}.$
+$f_{1}(0)=\int_{0}^{\tau}f_{\circ}(0,t)dt,~~~f_{2}(0)=\frac{f_{\circ}(0,0)}{f_{1}(0)}=\frac{f_{\circ}(0,0)}{\int_{0}^{\tau}f_{\circ}(0,t)dt}.$ (64)
 Therefore, using assumption 5.7, if $f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1},\bm{\theta}_{2})=f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$ for any $0\leq z\leq y\leq\tau$ and $\bm{x}\in\mathcal{X}$, then $(\bm{\theta}_{1},\bm{\theta}_{2})=(\bm{\theta}_{1}^{\prime},\bm{\theta}_{2}^{\prime})$.